### v1

# Large Language Models as Natural Selector for Embodied Soft Robot Design

Abstract
Designing soft robots is a complex and iterative process that demands cross-disciplinary expertise in materials science, mechanics, and control, often relying on intuition and extensive experimentation. While Large Language Models (LLMs) have demonstrated impressive reasoning abilities, their capacity to learn and apply embodied design principles—crucial for creating functional robotic systems—remains largely unexplored. This paper introduces RoboCrafter-QA, a novel benchmark to evaluate whether LLMs can learn representations of soft robot designs that effectively bridge the gap between high-level task descriptions and low-level morphological and material choices. RoboCrafter-QA leverages the EvoGym simulator to generate a diverse set of soft robot design challenges, spanning robotic locomotion, manipulation, and balancing tasks. Our experiments with state-of-the-art multi-modal LLMs reveal that while these models exhibit promising capabilities in learning design representations, they struggle with fine-grained distinctions between designs with subtle performance differences. We further demonstrate the practical utility of LLMs for robot design initialization. Our code and benchmark will be available to encourage the community to foster this exciting research direction1.
I Introduction
Soft robots offer distinct advantages over traditional rigid-bodied systems, particularly in complex, unstructured, and dynamic environments, where their inherent compliance enables safer and more adaptable interactions [1]. The compliance, adaptability, and distributed actuation of soft robots unlock new possibilities in applications ranging from minimally invasive surgery to flexible manufacturing and exploration [2, 3, 4]. However, these same characteristics introduce significant design challenges for soft robots. Unlike rigid robots with well-defined kinematic chains, soft robots possess virtually infinite degrees of freedom, exhibit non-linear material properties, and necessitate intricate coordination of morphology, actuation, and control. This complexity makes soft robot design a highly challenging multidisciplinary problem, traditionally relying on expert intuition, iterative prototyping, and computationally expensive simulations.
Figure 1 illustrates a conceptual shift in the paradigms of creature and robot design. While biological evolution (Fig. 1a) and human-engineered design (Fig. 1b) have driven progress, they are inherently slow and limited by human cognitive capacity. The emergence of Artificial Intelligence, particularly Large Language Models (LLMs), presents a new opportunity: LLMs as a form of ‘natural selector’ for robot design (Fig. 1c). LLMs, trained on vast datasets that encompass text and code, have demonstrated impressive capabilities in the understanding, generation, and even reasoning of natural languages. As discuss in [5], LLMs have the potential to revolutionize the robotic design process by serving as intelligent co-designers, however, there is no work that enables comprehensive and quantitative analyses and understanding of LLM for robot design. This raises a fundamental question: Can LLMs leverage their latent representations for the design of embodied soft robots, bridging the gap between high-level task descriptions and low-level morphological choices?
Figure 1: Conceptual comparison of creature/robot evolution paradigms. (a) Evolution-driven emergence of creatures in nature through selective pressures; (b) Traditional human-engineered robot design guided by intuition and expertise; (c) AI-driven robot design, where large language models (LLMs) act as a ‘natural selectorin the evolution of soft robotics, e.g., modular robot shown in the figure. This shift highlights the transition from biological evolution to human-driven engineering and finally to AI-mediated selection.
Existing LLM benchmarks focus primarily on general-purpose language understanding, code generation, or visual reasoning in non-embodied contexts. Although these benchmarks are valuable, they do not adequately assess the ability of LLMs to reason about physical systems, material properties, and the complex interplay between a robot’s morphology, its environment, and the task it must perform. In essence, current LLM benchmarks do not test the embodied design intelligence required for successful soft robot creation. The crucial question of whether LLMs can effectively learn and apply design principles for embodied soft robots remains largely unanswered.
To address this critical gap, we introduce RoboCrafter-QA, a novel benchmark specifically designed to evaluate the understanding of embodied soft robot design of multimodal LLMs. The entire pipeline, encompassing data generation, question formulation, and LLMs evaluation, is illustrated in Figure 2. RoboCrafter-QA leverages the EvoGym robotic simulator [6], a high-fidelity, computationally efficient platform for evolutionary robotics research. EvoGym allows us to generate a large-scale, controlled dataset of (environment, task, soft robot design, performance) tuples, providing a rich ground truth for evaluating soft robot design understanding. The core of RoboCrafter-QA is a challenging Question-Answering (QA) task: given a textual description of a simulated environment and task, a LLM must select the better-performing soft robot design from a pair of candidates. This task directly probes the LLM’s ability to learn and apply representations of embodied design principles, requiring it to reason about the intricate interplay of environmental factors, task objectives, and the morphological and material properties of the robot designs. Success hinges not on simple pattern matching, but on a deeper understanding of how design choices impact performance in a physically realistic simulation.
Figure 2: RobotCrafter-QA data generation and LLM evaluation pipeline.
Our benchmarking results, using state-of-the-art LLMs, reveal both the potential and the limitations of current models in this challenging domain. We find that while LLMs can learn to distinguish between significantly different designs, they struggle with fine-grained distinctions, highlighting the need for further advancements in representation learning for embodied design. Finally, beyond leveraging LLMs as robot design selectors, we show that they can also aid in refining and optimizing initial designs by incorporating task-specific constraints and physical reasoning.
This paper makes the following key contributions:
  1. 1.
    We introduce RoboCrafter-QA, a novel benchmark for evaluating the embodied soft robot design understanding of multi-modal LLMs, focusing on the evaluation of design representations.
  2. 2.
    We propose a challenging Question-Answering task that directly assesses LLMs’ ability to reason about the interplay between task requirements, environment constraints, and soft robot design, going beyond traditional pattern recognition.
  3. 3.
    We provide a comprehensive evaluation of state-of-the-art multi-modal LLMs on the RoboCrafter-QA benchmark, offering insights into their capabilities and limitations in applying embodied design principles.
  4. 4.
    We showcase that LLMs generate more effective robot design initializations than random initializations when provided with a textual task description.
II Related Work
II-A Soft Modular Robot Design and Optimization
Soft modular robots exploit distributed actuation and compliance to enhance flexibility and adaptability in dynamic environments [7, 8, 9, 10]. Unlike rigid robots with fixed kinematics, these systems dynamically alter morphology and stiffness using soft actuation mechanisms such as pneumatic actuators, shape memory alloy, and electroactive polymers. Recent studies have emphasized reconfigurability as a key advantage. Zhang et al. [11] introduced a lattice-based framework for morphology adaptation under external forces, while Li and Banerjee [12] developed modular designs for flexible attachment and detachment. Further approaches integrate sensor feedback with adaptive materials to enable real-time reconfiguration [13, 14]. However, optimizing soft modular robots remains challenging due to nonlinear deformations and complex material interactions. Traditional techniques, such as evolutionary algorithms and reinforcement learning [15, 16], provide heuristic-based solutions but struggle with scalability. Recent advances leverage generative AI for automated morphology discovery, with DiffuseBot [17] integrating diffusion-based generative modeling and differentiable physics, while Garcia et al. [18] combine data-driven and physics-based constraints for enhanced design efficiency.
Our work advances this field by benchmarking LLMs for automated robot design and proposing methods to systematically probe their reasoning in morphology optimization. Leveraging the EvoGym simulator as a high-fidelity testbed, we enable structured evaluations of LLM-driven design selection and optimization under realistic physics constraints.
II-B LLM Evaluation Benchmarks
General-purpose benchmarks such as BIG-Bench [19], MMLU [20], and C-Eval [21] assess LLM performance on linguistic fluency, knowledge retrieval, and reasoning. However, the rapid evolution of LLMs has led to specialized evaluations like ScienceQA [22, 23, 24] and MMLU-pro [25, 26], which test domain-specific reasoning in scientific and professional contexts. These benchmarks provide deeper insights into specialized knowledge but remain primarily focused on textual understanding. While these benchmarks evaluate language comprehension, they often overlook embodied intelligence—where physical constraints, environmental interactions, and morphology influence task success. Recent research has emphasized the need for evaluations that integrate physical reasoning, as LLMs often struggle with robotics-related decision-making [27]. Studies highlight the importance of specialized assessments that bridge high-level planning with real-world constraints [28, 29]. Without such evaluations, LLMs may demonstrate strong linguistic capabilities but fail to generalize effectively in real-world applications requiring physical interactions.
Existing benchmarks focus on textual reasoning but fail to measure how well LLMs incorporate task goals, environmental conditions, and morphological factors in robotics. To address this gap, our proposed RoboCrafter-QA benchmark provides a structured framework to assess how effectively LLMs reason about soft robot design choices. By incorporating multimodal constraints and physical reasoning, RoboCrafter-QA extends beyond conventional textual benchmarks, offering a more comprehensive evaluation of embodied intelligence in AI systems. Furthermore, it facilitates a deeper understanding of how LLMs navigate the intersection of design trade-offs, material properties, and task-specific constraints in robotics applications.
III RobotCrafter-QA Benchmark
III-A Benchmark Overview
The RoboCrafter-QA benchmark is designed to rigorously evaluate the embodied design reasoning capabilities of LLMs in the context of soft robot design over 12 tasks as illustrated in Figure 3. It comprises a curated dataset of Question-Answer (QA) pairs, $\{(Q_{i},G_{i})\}_{i=1}^{N}$, generated using the EvoGym simulator. Each question $Q_{i}=(D_{Q_{i}},R_{1_{i}},R_{2_{i}})$ presents a LLM with a textual prompt $D_{Q_{i}}$ and two candidate robot designs, $R_{1_{i}}$ and $R_{2_{i}}$, represented as voxel-based material matrices. The ground truth answer, $G_{i}\in\{A,B\}$, indicates the robot design that achieves the higher reward, $P(R,\pi,E)$, in the simulated environment $E$. The benchmark encompasses a diverse range of task environments, $E=(D,T,C)$, including locomotion, object manipulation, and climbing, each characterized by a textual description $D$, a task objective $T$, and environmental constraints $C$. An example of a question $Q_{i}$ and its components is illustrated in Figure 4. The difficulty of each question is modulated by the difference in performance rewards, $|P(R_{1_{i}},\pi,E)-P(R_{2_{i}},\pi,E)|$, between the candidate robot designs, enabling a nuanced assessment of LLMs’ ability to discern subtle design trade-offs. An example of various robot designs and their corresponding rewards for a pushing task and a bridge walking task is shown in Figure 5. The dataset is partitioned into Easy and Hard subsets based on this performance difference, providing a comprehensive evaluation of LLMs’ embodied design reasoning across varied complexity.
Figure 3: Robot design tasks in our RoboCrafter-QA benchmark.
Figure 4: Example RoboCrafter-QA question and LLM responses. This figure illustrates a sample question from the benchmark, showing the two robot designs (color-coded from input matrices), the question prompt, and the LLM responses with the ground truth. The simulated locomotion of the robot designs are illustrated at the bottom.
Figure 5: Example of robot designs with varied rewards.
III-B Data Generation
Design Space. Robot morphologies are evolved within a $5\times 5$ voxel-based design space, where each voxel denotes a material type: empty (0), rigid (1), soft (2), horizontal actuator (3), or vertical actuator (4). Structural integrity is enforced by maintaining fully connected components. Thus, a robot design $R$ is represented by a material matrix $M\in\{0,1,2,3,4\}^{5\times 5}$.
Evolution. The evolutionary process begins with an initial population of 30 unique robot designs, generated randomly. Each robot $R$ is evaluated in a designated Evogym environment $E$ using a Proximal Policy Optimization (PPO)-trained [30] controller $\pi(R,E)$, which controls the action signals applied on the actuator voxels. The controller is trained in the Evogym simulation to optimize the performance of the given robot structure. The fitness metric is the best reward, $P(R,\pi,E)$, reflecting per-task performance. We refer the reader to [6] for a detailed description of the parameter settings, including voxel specifications and environmental conditions.
Selection and Mutation. The top 50% of robots per generation are retained, while the remaining are replaced via mutation-based offspring generation. Mutations modify voxel properties at a low probability (about 10%), ensuring design diversity. Novelty constraints are enforced before acceptance to maintain uniqueness.
Termination and Data Storage. The evolutionary process continues with a constant population size of 30 until 3000 robots have been evaluated or an alternative termination criterion is met. The final population, along with its morphological data $M$ and performance data $P$, is archived for further analysis.
III-C Question Generation
Choice Construction. An automated pipeline transforms experimental data into structured multi-choice questions for LLM evaluation. Robot records are sorted by reward $P$ and partitioned into pairs, forming questions that require models to identify the higher-performing structure.
Context Integration. Task descriptions $T$ and difficulty levels are retrieved from predefined dictionaries. Each question includes environment-specific objectives, such as ”Climb as high as possible on a vertical wall” or ”Walk as far as possible on flat terrain” to provide necessary context $D_{Q_{i}}$.
Structured Representation. Each question $Q_{i}$ consists of:
  • A $5\times 5$ voxel grid representation $M$ detailing material types (rigid, soft, actuator, or empty).
  • Information on the actuation mechanism $\pi$ influencing movement.
  • A letter $G_{i}\in\{A,B\}$ indicating the superior robot based on empirical performance $P$.
III-D Benchmark Statistics
Table I provides a detailed overview of the RoboCrafter-QA benchmark dataset, showcasing its breadth and controlled difficulty across twelve distinct EvoGym tasks. Comprising a substantial 36,000 questions in total, with 3,000 questions dedicated to each environment, the benchmark ensures a robust evaluation. The ’Reward Diff’ column highlights a critical aspect of question difficulty, presenting the range of performance score differences between the paired robot designs. This range varies across environments, from subtle distinctions (e.g., Climber-v0 with 0.01 to 0.19) to more pronounced performance gaps (e.g., Jumper-v0 with 0.13 to 8.07), effectively modulating the challenge for LLMs. Furthermore, the ’Max Reward’ and ’Min Reward’ columns contextualize the performance landscape within these environments, demonstrating the variability in achievable and poor robot designs that contribute to the question generation process. This statistical breakdown underscores the dataset’s comprehensive nature, designed to rigorously challenge and differentiate the embodied design reasoning capabilities of multi-modal LLMs across a spectrum of task complexities and performance disparities.
IV Experiments
IV-A Evaluation Metrics
Accuracy: Measures the proportion of model-generated responses that match the expected answers. Given $N$ test queries, accuracy $A$ is computed as:
$A=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(O_{i}=G_{i})$ (1)
where $O_{i}$ is the model’s output, $G_{i}$ is the ground truth, and $\mathbb{I}(\cdot)$ is an indicator function returning 1 if responses match, 0 otherwise.
Consistency: Evaluates the stability of a large language model by measuring the reproducibility of the response in multiple trials. Each query is presented three times, and the consistency score $\kappa$ is defined as:
$\kappa=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbb{I}(O_{i}^{(1)}=O_{i}^{(2)})+\mathbb{I}(O_{i}^{(1)}=O_{i}^{(3)})}{2}$ (2)
where $O_{i}^{(1)},O_{i}^{(2)},O_{i}^{(3)}$ are outputs from three independent trials. Higher consistency indicates a more stable and deterministic model behavior.
TABLE I: Statistics of question sets across different environments, including the number of questions, reward difference range, maximum reward, and minimum reward.
Task Name # Questions Reward Diff Max Reward Min Reward
Pusher-v1 3000 0.04 to 1.39 0.21 -1.33
Jumper-v0 0.13 to 8.07 0.04 -8.44
Carrier-v1 0.90 to 6.62 3.60 -5.85
Walker-v0 0.74 to 2.89 4.09 -1.21
UpStepper-v0 0.08 to 2.51 1.64 -1.27
Balancer-v0 0.12 to 1.05 0.13 -1.41
Climber-v0 0.01 to 0.19 0.16 -0.18
Pusher-v0 0.49 to 3.07 3.07 -1.66
Carrier-v0 0.53 to 3.50 2.81 -2.86
GapJumper-v0 0.08 to 2.62 2.28 -0.77
BridgeWalker-v0 0.39 to 3.69 4.04 -1.49
BidirectionalWalker-v0 0.31 to 4.16 1.81 -3.69
Summary 36000 0.01 to 8.07 4.09 -8.44
Difficulty-Weighted Accuracy (DWA): Quantifies the model’s ability to discern subtle differences in robot task performance, with higher weights assigned to harder questions (smaller reward differences). Let $N_{tasks}$ be the total number of robot tasks. For each robot task $t\in\{1,...,N_{tasks}\}$, define:
$\text{Difficulty}_{t,i}=1-\frac{\Delta P_{t,i}-\Delta P_{t}^{\text{min}}}{\Delta P_{t}^{\text{max}}-\Delta P_{t}^{\text{min}}}$ (3)
$\text{DWA}_{t}=\frac{\sum_{i=1}^{N_{t}}\mathbb{I}(\text{Correct}_{t,i})\times\text{Difficulty}_{t,i}}{\sum_{i=1}^{N_{t}}\text{Difficulty}_{t,i}}$ (4)
Cross-Task Normalization: Aggregates task-specific DWAs while eliminating scale variance:
Global $\text{Global DWA}=\frac{1}{N_{tasks}}\sum_{t=1}^{N_{tasks}}\text{DWA}_{t}$ (5)
where $\Delta P_{t,i}=P_{t,i}^{\text{correct}}-P_{t,i}^{\text{incorrect}}$ is the reward gap between correct and incorrect options for question $i$ in robot task $t$, $\Delta P_{t}^{\text{min}}$ and $\Delta P_{t}^{\text{max}}$ are the minimum and maximum reward gaps within task $t$, $\mathbb{I}(\cdot)$ is the indicator function, and $N_{t}$ is the number of questions per task. Higher DWA indicates better performance on discriminative tasks with small reward differences.
IV-B Benchmarking Results of Robot Design Selection
We conduct a comparative evaluation of four most advanced available Large Language Models (LLMs) across 12 diverse environments, using metrics categorized into ‘Easy’ and ‘Hard’ tasks. The summarized results of accuracy and consistency are presented in Table II.
TABLE II: Accuracy and Consistency metrics across different LLMs for various robotic design tasks.
Category Task Name GPT-o3-mini Gemini-2.0-flash Gemini-1.5-Pro Grok-2
Accuracy Consistency Accuracy Consistency Accuracy Consistency Accuracy Consistency
Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard
Locomotion Walker-v0 57.65 56.58 73.11 72.73 71.40 63.00 94.15 94.75 75.40 65.20 96.10 96.75 71.33 61.70 90.60 89.85
BridgeWalker-v0 49.65 44.86 72.17 77.99 75.40 61.80 95.70 91.55 81.57 68.20 97.95 97.50 78.40 63.67 91.20 89.25
BidirectionalWalker-v0 51.46 52.90 72.91 73.26 55.23 56.11 93.60 95.31 53.97 55.03 97.15 97.00 55.30 53.50 88.20 90.70
UpStepper-v0 63.57 53.72 79.02 78.45 72.97 56.70 95.95 93.95 72.97 61.10 97.00 95.40 74.00 58.43 94.85 92.00
GapJumper-v0 58.88 53.63 77.62 76.31 71.53 58.83 94.25 94.85 78.97 70.00 97.95 96.90 74.57 62.93 92.75 90.70
Average 56.24 52.34 74.97 75.75 69.31 59.29 94.73 94.08 72.58 63.91 97.23 96.71 70.72 60.05 91.52 90.50
Object Manipulation Carrier-v0 60.60 57.17 78.58 81.99 73.70 61.57 96.65 94.85 78.43 63.43 97.30 96.30 75.70 63.20 94.90 91.80
Carrier-v1 55.79 51.46 72.06 68.75 72.30 61.03 95.00 92.80 72.07 62.10 96.45 95.60 75.37 65.67 91.40 89.95
Pusher-v0 52.07 53.73 72.63 73.91 68.27 61.83 94.65 93.80 78.77 64.53 97.50 96.40 72.23 62.43 90.30 89.35
Pusher-v1 59.26 59.34 76.37 76.36 62.70 62.18 95.05 95.87 65.33 61.60 96.70 97.75 61.73 59.43 93.95 93.15
Average 56.93 55.43 74.91 75.25 69.24 61.65 95.34 94.33 73.65 62.92 96.99 96.51 71.23 62.68 92.64 91.06
Climbing & Balancing Climber-v0 54.59 53.33 85.29 85.67 58.83 58.60 96.15 95.90 59.47 58.10 97.70 97.00 58.57 58.63 92.35 93.85
Jumper-v0 9.70 33.42 91.36 81.32 47.97 50.70 93.25 91.99 39.83 49.00 96.20 95.95 27.27 41.90 93.20 90.90
Balancer-v0 63.38 60.96 78.72 79.46 69.10 68.26 96.45 95.66 68.27 71.47 97.40 98.20 71.97 69.17 91.85 92.80
Average 42.56 49.24 85.12 82.15 58.63 59.19 95.28 94.52 55.86 59.52 97.10 97.05 52.60 56.57 92.47 92.52
Average 55.34 52.88 79.28 77.58 66.62 60.05 95.07 94.27 68.75 62.48 97.12 96.73 66.37 60.06 92.13 91.19
Finding 1: Performance Hierarchy Among Models. The performance of the evaluated models exhibits a clear hierarchy across both accuracy and consistency metrics. Gemini-1.5-Pro demonstrates the highest accuracy in both easy tasks (68.75%) and hard tasks (62.48%), surpassing other models across various environments. This suggests that Gemini-1.5-Pro is more effective in identifying optimal robot designs across a range of task complexities. In contrast, Gemini-2.0-flash and Grok-2 exhibit moderate performance, with average easy accuracies of 66.62% and 66.37%, respectively, and hard accuracies around 60%, showing competitive but slightly lower capabilities compared to Gemini-1.5-Pro. Conversely, GPT-o3-mini exhibits the lowest accuracy (55.34% in easy tasks and 52.88% in hard tasks), with notable struggles in tasks like Jumper-v0 (9.70% easy), indicating challenges in generalizing effectively, particularly in complex scenarios requiring nuanced decision-making.
Finding 2: Trade-off Between Accuracy and Consistency. In this evaluation, there is no clear trade-off between accuracy and consistency; instead, performance in both metrics aligns with overall model capability. Gemini-1.5-Pro achieves the highest consistency (97.12% in easy tasks and 96.73% in hard tasks) alongside its top accuracy, highlighting its ability to maintain stable and reliable predictions across all task types. Gemini-2.0-flash follows with strong consistency (95.07% easy, 94.27% hard), outperforming Grok-2 (92.13% easy, 91.19% hard) despite similar accuracy levels, making it suitable for tasks where repeatability is critical. In contrast, GPT-o3-mini shows significantly lower consistency (79.28% easy, 77.58% hard) and accuracy, suggesting that its predictions are less stable and less reliable, particularly as task complexity increases.
Finding 3: Sensitivity to Task Difficulty. The performance of the models reveals sensitivity to task difficulty, with all models showing reduced precision in harder tasks where distinguishing between robot designs may be more challenging. For example, in Walker-v0, the precision drops from 75.40% (easy) to 65.20% (hard) for Gemini-1.5-Pro and from 57.65% (easy) to 56.58% (hard) for GPT-o3-mini, indicating that subtler performance differences in harder tasks pose difficulties. This trend is pronounced in tasks such as Jumper-v0, where the accuracy of GPT-o3-mini increases from 9. 70% (easy) to 33. 42% (hard), but remains low, suggesting an inconsistent sensitivity to task complexity. Gemini-1.5-Pro maintains relatively high performance across both easy and hard tasks, while Gemini-2.0-flash sustains strong consistency (e.g., 94.15% easy and 94.75% hard in Walker-v0), implying it relies on stable predictions even when accuracy decreases slightly. GPT-o3-mini struggles most, with low accuracy and consistency in complex tasks, highlighting its limitations in fine-grained evaluation.
Finding 4: Weaknesses of Models in Certain Environments The tasks Jumper-v0 and BidirectionalWalker-v0 reveal significant weaknesses in the evaluated model. All models exhibit very low accuracy, but high consistency, indicating a persistent inability to identify effective designs. This may be attributed to the specific characteristics and descriptions of the environments and tasks. For example, Jumper-v0 likely demands precise timing or control (e.g., jumping mechanics), which LLMs may not be adept at handling due to their lack of specialized reasoning for dynamic tasks. The bidirectional nature of the task BidirectionalWalker-v0 may also introduce ambiguous or conflicting cues (e.g., deciding between forward and backward movement), confusing the models about the correct course of action. The models’ architectures might not be suited to handle bidirectional decision-making, which requires understanding symmetric or reversible actions, leading to errors in design selection.
TABLE III: Global DWA metrics across different LLMs.
Environment GPT-o3-mini Gemini-2.0-Flash Gemini-1.5-Pro Grok-2
Walker-v0 58.68 65.52 67.82 64.71
BridgeWalker-v0 45.39 66.50 72.33 68.11
BidirectionalWalker-v0 52.11 55.55 54.31 54.58
UpStepper-v0 58.54 63.23 65.46 64.74
GapJumper-v0 57.11 64.38 74.10 68.94
Carrier-v0 60.07 66.92 69.80 69.08
Carrier-v1 53.84 66.02 66.55 70.65
Pusher-v0 53.01 63.81 68.77 65.80
Pusher-v1 60.53 62.19 63.69 60.77
Climber-v0 54.12 58.26 58.00 57.69
Jumper-v0 19.31 49.19 44.69 34.86
Balancer-v0 62.70 69.28 70.95 71.39
Average 52.96 61.99 63.72 61.66
Figure 6: Error distribution across difficulty levels for LLMs.
IV-C Performance Analysis Across Reward Difference Level
To assess the ability of LLMs to select better designs across varying levels of difficulty, we use the DWA metric, as introduced in Section IV-A. The results for these metrics are presented in Table III. Table III presents a comparative analysis of global Difficulty-Weighted Accuracy (DWA) metrics across four distinct LLMs: GPT-o3-mini, Gemini-2.0-Flash, Gemini-1.5-Pro, and Grok-2. The table breaks down the DWA performance for each LLM across twelve different EvoGym environments, providing a granular view of their design selection capabilities under varying task contexts. Notably, the DWA scores fluctuate across environments and LLMs, indicating varying degrees of success in discerning superior robot designs, weighted by the difficulty of each question. The ’Global Average’ row summarizes the overall performance of each model across all environments, offering a consolidated metric for comparison. Gemini-1.5-Pro emerges with the highest global average DWA of 63.72%, suggesting a marginally superior capability in embodied design reasoning within the RoboCrafter-QA benchmark compared to the other evaluated models.
Furthermore, we visualize the distribution of incorrect answers in different levels of reward difference in Figure 6. In particular, a significant proportion of errors made by LLM occur within the 0.8–1.0 difficulty interval, further highlighting their current limitations in making fine-grained distinctions between designs.
IV-D Ablation Study on Minor Prompt Design
We conduct an ablation study on prompt design to determine the key factors influencing the LLMs’ ability to make correct selections and assess how different information affects their performance. Given that the numerical representation of voxel materials is crucial for design, we focus our ablation study on task and actuator descriptions. Additionally, we perform an experiment where we modify the prompt to instruct the LLMs to select the Worse design instead of the better one, further analyzing the robustness of decision-making process of LLMs.
TABLE IV: Ablation study results on prompt design
GPT-o3-mini Gemini-2.0-flash Gemini-1.5-pro Grok-2
NoAct 54.09 66.93 69.78 65.82
NoEnv 52.08 62.47 62.50 62.91
Worse 51.93 58.48 67.87 61.24
Full 55.34 66.62 69.75 66.37
The ablation study on prompt design highlights the critical role of task and actuator descriptions in enabling language models (LLMs) to select optimal designs. The results in Table IV show that removing the task description (NoEnv) significantly reduces performance across all models, for example, from 55.34% to 52.08% for GPT-o3-mini and from 69.75% to 62.50% for Gemini-1.5-pro, underscoring its importance in guiding the decision-making process of LLMs. In contrast, omitting the actuator description (NoAct) has a minimal effect, with performance remaining stable or slightly varying. This may indicate that without actuator information, LLMs assume that the actuator is capable of maximizing the design’s reward. Additionally, when instructed to select the Worse design, the models exhibit lower accuracy compared to the full-information prompt (e.g., 66.62% to 58.45% for Gemini-2.0-flash), suggesting they are less adept at identifying inferior designs, likely due to a training bias toward selecting better ones. These findings emphasize the need for comprehensive task descriptions to maximize LLM performance in design selection tasks.
TABLE V: Comparison of design strategies.
Method Valid Design Reward Mean Reward Variance
Random 50/50 0.044 0.405
w/o Reference Knowledge 19/50 -0.607 0.255
w/ Reference Knowledge 50/50 0.115 0.035
IV-E Case Study of Robot Design initialization via LLM
In addition to selecting the optimal design in multiple-choice tasks, we aim to evaluate whether current large language models (LLMs) exhibit zero-shot capabilities for design initialization using reference information. Specifically, we use Walker-v0 and Jumper-v0 as reference environments, providing the LLM with 10 high-reward and 10 low-reward designs from each. The model is then tasked with generating a feasible structure for the GapJumper-v0 environment, leveraging environment details, $E=(D,T,C)$, along with the design space. We select Gemini-1.5-Pro as the best-performing model based on results from Table II and compare its performance against randomly generated designs and those produced without reference knowledge from other environments. The results are summarized in Table V.
With reference knowledge, all 50 designs were valid, yielding the highest mean reward of 0.115 and the lowest variance of 0.035, reflecting strong and consistent performance. Without reference knowledge, only 19 of 50 designs were valid, with a mean reward of -0.607 and a variance of 0.255. The random baseline achieved 50 valid designs but had a lower mean reward of 0.044 and a higher variance of 0.405, indicating poorer and less stable outcomes. These results show the model effectively transfers knowledge from reference environments to initialize designs in a new environment, demonstrating zero-shot design generation capabilities.
V Conclusion
This paper introduces RoboCrafter-QA, a benchmark for evaluating multi-modal LLMs’ embodied design reasoning in soft robot design. Using the EvoGym simulator, we generated a diverse QA dataset capturing relationships between task requirements, environmental constraints, and robot morphologies. Experiments show that while LLMs perform well on simpler design selection, they struggle with subtle trade-offs and complex environments. RoboCrafter-QA provides a valuable tool for assessing and advancing LLMs’ design intelligence, enabling more automated and intelligent soft robot design.
Limitations and Future Work. This work has limitations suggesting future research directions. First, using pre-trained PPO policies could be improved by exploring LLM-driven control policy optimization. Second, the $5\times 5$ voxel grid restricts design complexity; a larger design space presents simulation challenges. Third, sim-to-real transfer, including material properties and controller transferability, is needed for real-world applications. Finally, multi-modal prompts (visual, tactile) could enhance LLMs’ embodied design understanding. Addressing these limitations will advance LLM-driven soft robot design.